A valószínűségszámítás és matematika statisztika gyakorlaton való részvétel

dátum:  IX. 7. IX. 14. IX. 21. IX. 28. X. 5. X. 12. XI. 2.

diákok:

Bordi István  +  +  +  +  +
Boros László  +  +  −  −  −
Böszörményi Zsolt  +  +  +  +
Csipkés László  +  +  +  +  −
Ferenczik Zoltán  +  +  −  +  +
Fogas János  +  +  +  +  −
Havasi Balázs  +  +  +  +  −
Hennel Zoltán  +  −  +  −
Illés József  +  +  +  +  +
Jagó Gábor Jácint  +  +  +  +  +
Kiss Róbert  +  +  +  +  +
Lászlók Zsolt  +  +  +  +  −
Mészáros Gábor  +  −  +  −  −
Mezei Tamás  +  +  +  +  +
Lázár Gábor  +
Molnár Adrienn  +  +  +  +
Molnár Tibor  +  +  −  −
Molnár Zoltán  +  +  +  +  +
Nagy András  +  +  +  +  +
Paper Andrea  +  +  +  −
Pelejtei Gergely  +  +  −  +  +
Szalontai Gábor  +  +  +  −  −
Tordai Jácint  +  +  +  +  −
Ujjobágy Attila  +  +  −  +  +
Vágner József  +  −  +  −  +
Virág Gabriella  +  +  +  +  +
Zádor István  +  +  +  +  +